# SVD Based Least Squares for X-Ray Pneumonia Classification Using Deep Features

### v1

@@ 3.3 Baselines with Pretraining and Fine Tuning Using Gradient Descent
By using the gradient descent, we train the following models on our pneumonia classification dataset: 1) ResNet-18 trained from scratch, 2) ResNet-18 where all parameters trained after transfer learning initialization, 3) ResNet-18 where the last linear layer is fine-tuned after transfer learning initialization and 4) DINO ViT-B/16 model fine-tuned using a linear classifier. The results of these baseline models are presented in Table 1.
@@ 3.4 Training Specifications
In the training of each of the gradient descent based methods, we used the Adam [27] optimizer with weight decay. We performed a grid-search to find the optimal values for the hyperparameters, and k-fold cross validation with $k=5$ on each grid-search combination. The training parameters found are presented in Table 2. As we have a fixed test set, for each k-fold we divided the training set into 80% (training) and 20% (validation) portions and reported the test set results of the model with the best validation F1-score. Furthermore, for the SVD-LS classification framework, we obtain a linear classifier on top of the following models: 1) ResNet-18, 2) ResNet-50, and 3) DINO ViT-B/16. Extracted features from these models are used as inputs to our SVD-LS classifier. This allows us to leverage the representation power of deep networks while retaining the efficiency by closed-form solutions. We again perform grid-search and k-fold cross-validation with $k=5$ to determine the optimal number of principal components to retain after SVD-based dimensionality reduction.
Model Epoch Optimizer Learning Rate Weight Decay Batch Size Loss Scheduler
ResNet-18 From Scratch 30 Adam 0.0001 1e-05 16 Cross Entropy StepLR
ResNet-18 Transfer Learn 30 Adam 0.0005 1e-05 32 Cross Entropy StepLR
ResNet-18 Fine-Tuning 30 Adam 0.001 1e-05 16 Cross Entropy StepLR
DINO Fine-Tuning 30 Adam 0.001 1e-04 16 Cross Entropy StepLR
Table 2: Training details of gradient descent based models.
@@ 4 Results
The results of our experiments, shown in Table 1, demonstrate clear differences in performance across various models for multi-class pneumonia classification on chest X-ray images. Among the evaluated models, the DINO SVD-LS model, which combines a self-supervised transformer with SVD-based Least Squares, outperformed all other models on the test set, achieving an F1-score of 84.27% and an accuracy of 90.28%. This highlights the superior generalization capabilities of self-supervised models using SVD-LS. ResNet-18 fine-tuning (M3), which involves training only the last layer, demonstrated the second-best test performance with an F1-score of 77.82% and an accuracy of 86.00%. While ResNet-18 SVD-LS (M4) and ResNet-50 SVD-LS (M5) also utilized SVD-LS, their performance was inferior to DINO SVD-LS, with M5 underperforming, indicating higher model complexity without proper feature extraction may hurt performance.
To better understand the quality of learned features, we visualize the t-SNE [28] projections of feature vectors from both the DINO ViT-B/16 and ResNet-18 transfer-learned models (see Figure 2). While both show some class separation, DINO features form more compact and well-separated clusters, indicating stronger class-discriminative structure. This makes DINO features particularly suitable for linear classifiers. Despite their separability, these high-dimensional features may still contain redundant or noisy components. SVD addresses this by projecting them onto a lower-dimensional subspace spanned by the most informative directions, reducing overfitting and improving numerical stability. The t-SNE plots qualitatively suggest a low-dimensional manifold structure, which SVD can capture effectively, enabling the least squares classifier to generalize better in this reduced space.
To validate performance differences between each model, we applied the McNemar-Bowker test [29], a statistical method for comparing paired classifier predictions. This test examines the disagreements between models using a contingency table and determines statistical significance through a chi-squared distribution. As seen in Table 3, the small confidence values confirm the statistical differences in model outputs.
5 Conclusion
In this work, we introduced an alternative classification framework for multi-class pneumonia classification using chest X-ray images, based on Singular Value Decomposition-based Least Squares (SVD-LS). Using feature representations from state-of-the-art self-supervised and transfer learning models, our approach achieves a compelling balance between accuracy and efficiency. Notably, our results demonstrate that combining SVD-LS with self-supervised features from a DINO ViT-B/16 transformer model yields the highest test performance among all evaluated models and baselines.
Unlike conventional deep learning pipelines that rely on iterative optimization, our method provides a closed-form solution with significantly reduced computational cost and training time, making it particularly suitable for real-time applications and deployment in resource-constrained clinical environments. In future work, we plan to explore the integration of uncertainty estimation techniques into the SVD-LS pipeline to enhance its reliability in clinical decision support.
@@ 5 Conclusion
M7 0.000 0.000 0.001 0.074 0.001 $0$
Table 3: Pairwise McNemar-Bowker Test [29] results comparing classifier predictions. Significant p-values (<0.05) indicate meaningful performance differences.

### v2

@@ 3.4 Baselines with Pretraining and Fine Tuning Using Gradient Descent
By using gradient descent, we trained the following models on our pneumonia classification dataset: 1) ResNet-18 trained from scratch, 2) ResNet-18 where all parameters trained after transfer learning initialization, 3) ResNet-18 where the last linear layer is fine-tuned after transfer learning initialization and 4) DINO ViT-B/16 model fine-tuned using a linear classifier. The results of these baseline models are presented in Table 1.
@@ 3.5 Training Specifications
In the training of each of the gradient descent-based methods, we used the Adam [29] optimizer with weight decay. We performed a grid search to find the optimal values for the hyperparameters, and k-fold cross validation with $k=5$ for each grid-search combination. The training parameters we used can be found in our GitHub repository. As we have a fixed test set, for each k-fold we split the training set into 80% (training) and 20% (validation) portions and reported the test set results of the model that achieved the best validation F1-score. Furthermore, for the SVD-LS classification framework, we obtained a linear classifier on top of the following models: 1) ResNet-18, 2) ResNet-50, and 3) DINO ViT-B/16. Extracted features from these models were used as inputs to our SVD-LS classifier. This allowed us to leverage the representation power of deep networks while retaining the efficiency through closed-form solutions. We also performed grid search and k-fold cross-validation with $k=5$ to determine the optimal number of principal components to retain in SVD-LS.
@@ 4 Results
The results of our experiments, shown in Table 1, demonstrate clear differences in performance across various models for multi-class pneumonia classification on chest X-ray images. Among the evaluated models, the DINO SVD-LS model, which combines a self-supervised transformer with SVD-based least squares, outperformed all other models on the test set, achieving an F1-score of 84.27% and an accuracy of 90.28%. This highlights the superior generalization capabilities of self-supervised models using SVD-LS. ResNet-18 fine-tuning (M3), which involves training only the last layer, demonstrated the second-best test performance with an F1-score of 77.82% and an accuracy of 86.00%. While ResNet-18 SVD-LS (M4) and ResNet-50 SVD-LS (M5) used SVD-LS as well, their performance was inferior to DINO SVD-LS, suggesting that increased model complexity without corresponding feature quality may degrade performance.
To better understand the quality of learned features, we visualized the t-SNE [30] projections of feature vectors from both the DINO ViT-B/16 and ResNet-18 transfer-learned models (see Figure 2). While both show some class separation, DINO features form more compact and well-separated clusters, indicating stronger class-discriminative structure. This makes DINO features particularly suitable for linear classifiers. Despite their separability, these high-dimensional features may still contain redundant or noisy components. SVD addresses this by projecting them onto a lower-dimensional subspace spanned by the most informative directions, reducing overfitting and improving numerical stability. The t-SNE plots qualitatively suggest a low-dimensional manifold structure, which SVD can capture effectively, enabling the least squares classifier to generalize better in this reduced space.
To assess performance differences, we used the McNemar-Bowker test [31], which compares paired classifier predictions via a chi-squared test on a contingency table. As shown in Table 2, the low $p$-values (<0.05) indicate statistically significant differences between model outputs.
@@
M7 0.000 0.000 0.001 0.074 0.001 0.000 -
Table 2: Pairwise McNemar-Bowker test [31] results comparing the predictions of the models evaluated in Table 1.
5 Conclusion
In this work, we introduced an alternative classification framework for multi-class pneumonia classification using chest X-ray images, based on Singular Value Decomposition-based Least Squares (SVD-LS). Using feature representations from state-of-the-art self-supervised and transfer learning models, our approach achieves a compelling balance between accuracy and efficiency. In particular, our results demonstrated that SVD-LS combined with DINO ViT-B/16 self-supervised features yields the highest test performance among all models and baselines evaluated. Unlike conventional deep learning pipelines that rely on iterative optimization, our method provides a closed-form solution with significantly reduced computational cost and training time, making it particularly suitable for real-time applications and deployment in resource-constrained clinical environments. In future work, we plan to explore the integration of uncertainty estimation techniques into the SVD-LS pipeline to enhance its reliability in clinical decision support.